OpenStudy (anonymous):

Solve for x. 11x + 4 37

hartnn (hartnn):

subtract 4 from both sides, what u get?

hartnn (hartnn):

11x+4-4=37-4 can u solve further ?

OpenStudy (anonymous):

no don't know how

hartnn (hartnn):

what is 37-4 = ?

OpenStudy (anonymous):

33

hartnn (hartnn):

11x = 33 so whats x ?

OpenStudy (anonymous):

33

hartnn (hartnn):

11x = 33 x = 33/11 = 3

hartnn (hartnn):

ok?

OpenStudy (anonymous):

how you got 3

OpenStudy (amistre64):

11 times (something) equals 33